OpenStudy (anonymous):

MEDALS WILL BE GIVEN !!!!!!!!!!!!!!!!!1111! A line has a slope of 9/4 and a y-intercept of −14. Which equations represent the line? Choose all answers that are correct. A. -9x+4y = -56 B. 9x + 4y = 56 C. y= 9/4x -14 D. y = 9/4x + 14

OpenStudy (anonymous):

the standard slope-intercept form for an equation of a line is y+=+highlight_green%28m%29x+%2B+highlight%28b%29 where m is the slope and b the y-intercept. y = (9/4)x - 14 0r Standard Form: 9x-4y = 56

OpenStudy (anonymous):

@MarshallEatsFood does that help?

OpenStudy (anonymous):

not really....

OpenStudy (anonymous):

its asking for the letters not the solved question

OpenStudy (anonymous):

and even then i still dont understand it

OpenStudy (anonymous):

@kittenlover731

OpenStudy (anonymous):



OpenStudy (anonymous):

y = 9/4x - 14 <=== this one y = 9/4x - 14 --- multiply entire equation by common denominator of 4 4y = 9x - 56 --- subtract 9x from both sides -9x + 4y = -56 <=== this one 9x - 4y = 56 <== this would also be one, but it is not an answer choice so, A and C

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Yw :)

OpenStudy (anonymous):

wait could you explain how you did that but with more detail please??

OpenStudy (anonymous):

@Buttercup214

OpenStudy (anonymous):

The slope represents x and y intercept represents y, therefore the slope is y=9/4x -14

OpenStudy (anonymous):

A.-9x+4y=-56 is another equation

OpenStudy (anonymous):

still confused

OpenStudy (anonymous):

If you put -9x+4y=-56 in a slop intercept form it would be y=9/4x -14

OpenStudy (anonymous):

sorry i just dont get this ;-;

OpenStudy (anonymous):